# Xinlu Zhang1, Chenxin Tian2, Xianjun Yang1, Lichang Chen3, Zekun Li1, Linda Ruth Petzold 1

### v4

@@ Abstract
Instruction-finetuning (IFT) has become crucial in aligning Large Language Models (LLMs) with diverse human needs and has shown great potential in medical applications. However, previous studies mainly fine-tune LLMs on biomedical datasets with limited diversity, which often rely on benchmarks or narrow task scopes, and hence significantly limit the effectiveness on their medical instruction-following ability and generalizability. To bridge this gap, we propose creating a diverse, machine-generated medical IFT dataset, MedInstruct-52k, using GPT-4 and ChatGPT with a high-quality expert-curated seed set. We then fine-tune LLaMA-series models on the dataset to develop AlpaCare. Despite using a smaller domain-specific dataset than previous medical LLMs, AlpaCare not only demonstrates superior performance on medical applications, with up to 38.1% absolute gain over best baselines in medical free-form instruction evaluations, but also achieves 6.7% absolute gains averaged over multiple general domain benchmarks. Human evaluation further shows that AlpaCare consistently outperforms best baselines in terms of both correctness and helpfulness. We offer public access to our data, model, and codebase in https://github.com/XZhang97666/AlpaCare.
@@ 1 Introduction
Recent advancements in the training of large language models (LLMs) have placed a significant emphasis on instruction-finetuning (IFT), a critical step in enabling pre-trained LLMs to effectively follow instructions (Ouyang et al., 2022; Longpre et al., 2023; Wei et al., 2022). However, relying solely on NLP benchmarks to create instructional datasets can lead to ‘game-the-metric’ issues, often failing to meet actual user needs (Ouyang et al., 2022). To better align with human intent, Wang et al. (2023b) introduces the concept of fine-tuning LLMs using diverse machine-generated instruction-response pairs. Subsequent works further highlight the importance of diversity in IFT datasets (Taori et al., 2023; Xu et al., 2023a; Chiang et al., 2023; Chen et al., 2023; Bukharin & Zhao, 2024). However, how to improve dataset diversity in the medical domain for aligning with various user inquiries is still underexplored.
LLMs have demonstrated significant potential in the medical domain across various applications (Nori et al., 2023a; b; Singhal et al., 2022; 2023; Liévin et al., 2023; Zhang et al., 2023). To alleviate privacy concerns and manage costs, several medical open-source LLMs (Han et al., 2023; Li et al., 2023b; Wu et al., 2023; Xu et al., 2023b) have been developed by tuning LLaMA (Touvron et al., 2023a; b) on medical datasets. Even substantial volumes, these datasets are limited in task scopes and instructions, primarily focusing on medical benchmarks or specific topics, due to the high cost of collecting real-world instruction datasets (Wang et al., 2023b), particularly when extending further into the medical domain(Jin et al., 2021; 2019). This lack of diversity can negatively impact the models’ ability to follow instructions in various medical applications and their effectiveness in the general domain. Therefore, there is an urgent need for a method to generate diverse medical IFT datasets that align with various domain-specific user inquiries while balancing cost.
@@
To automatically generate a broader array of tasks for training, we prompt GPT-4 to create instructions for new medical tasks by leveraging the existing clinician-curated tasks as demonstrations. After generating tasks and conducting deduplications, we employ ChatGPT to provide responses to the valid tasks. Consequently, we compile a 52k medical self-instruct dataset, MedInstruct-52k, which supervises the tuning on the LLaMA series models (Touvron et al., 2023a; b), resulting in AlpaCare. Due to the limited number of test sets available for evaluating medical LLMs in terms of instruction-following ability and medical capacity, we introduce a new clinician-crafted free-form instruction evaluation test set, MedInstruct-test, covering medical tasks across different difficulty levels.
@@
  • To address the challenge of generating cost-effective, high-quality IFT data for LLM alignment in various medical applications, we propose a pipeline to create a diverse medical machine-generated IFT dataset for tuning LLMs.
@@
  • We release, MedInstruct-52K, a diverse medical IFT dataset comprising 52K instruction-response pairs and, MedInstruct-test, a test set of 216 clinician-crafted novel medical tasks, to facilitate the building and evaluation of medical LLMs.
@@ 2 Related Works
IFT. Closed-form IFT (Wei et al., 2022; Longpre et al., 2023) creates IFT datasets from existing NLP benchmarks to improve model generalization on new tasks. For instance, FLAN (Wei et al., 2022), T0 (Sanh et al., 2022), and Flan-T5 (Chung et al., 2022) construct their IFT datasets using existing NLP tasks with carefully designed instructions and demonstrate that fine-tuning with diverse task instructions enhances performance on unseen tasks. However, these closed-form instructions are often simpler compared to cases in real-world scenarios, making their models fail to align with various real-world user intentions. Alternatively, Ouyang et al. (2022) collects a diverse IFT dataset with real-world instructions and responses, which is rich in both instruction forms and task types, and trains GPT-3 Brown et al. (2020) to obtain InstructGPT (Ouyang et al., 2022) on this dataset, showing promising results in aligning with diverse actual user needs. Due to the closed-source propriety of strong LLMs (e.g. ChatGPT and GPT-4), various open-source instruction fine-tuned models (Taori et al., 2023; Xu et al., 2023a; Chiang et al., 2023; Peng et al., 2023) have been proposed to fine-tune open-source LLMs using datasets obtained from these strong teacher models to enhance their instruction-following abilities . Alpaca (Taori et al., 2023) creates a 52k diverse machine-generated IFT dataset by distilling knowledge from the "teacher" Text-Davinci-003 (Hinton et al., 2015; Li et al., 2022). Peng et al. (2023) utilizes the same instructions with Alpaca but adopts GPT-4 as the "teacher" LLM to generate higher-quality and more diverse responses to improve the model’s alignment on 3H (Helpfulness, Honesty, and Harmlessness) (Askell et al., 2021). Vicuna (Chiang et al., 2023) is trained on the ShareGPT data (Sharegpt, 2023), which contains actual ChatGPT users’ diverse instructions, obtaining strong response quality and instruction-following ability. However, creating diverse IFT datasets for aligning models with various domain-specific user intentions in the medical domain remains underexplored.
@@
LLMs in Biomedicine. Closed-source LLMs have demonstrated significant proficiency in the medical domain (OpenAI, 2022; 2023; Singhal et al., 2023; 2022). ChatGPT demonstrates promise in the US Medical Exam (Kung et al., 2023) and serves as a knowledge base for medical decision-making (Zhang et al., 2023). The MedPaLM (Singhal et al., 2022; 2023) have shown performance in answering medical questions on par with that of medical professionals. GPT-4 (OpenAI, 2023) obtains strong medical capacities without specialized training strategies on medical domain or engineering for solving clinical tasks Nori et al. (2023a; b). Due to privacy concerns and high costs, several open-source medical LLMs (Xu et al., 2023b; Han et al., 2023; Li et al., 2023b; Wu et al., 2023) have be built by tuning open-source base model, such as LLaMA (Touvron et al., 2023a; b), on medical corpus. However, due to the high cost of collecting diverse real-world user instructions (Wang et al., 2023b), their datasets are limited in diversity and primarily focus on medical benchmarks or narrow task scopes, such as doctor-patient conversations. ChatDoctor (Li et al., 2023b) is fine-tuned using 100k online doctor-patient dialogues, while Baize-Healthcare (Xu et al., 2023b) employs about 100k Quora and MedQuAD dialogues. MedAlpaca (Han et al., 2023) utilizes a 230k dataset of question-answer pairs and dialogues. PMC-LLAMA (Wu et al., 2023) continually trains LLaMA with millions of medical textbooks and papers, and then tunes it with a 202M-token dataset formed by benchmarks and dialogues in the IFT stage. However, these datasets are limited in diversity, mainly focusing on benchmarks or within certain topics, hampering models’ medical instruction-following ability and generalizability. To address this, we propose building a cost-effective diverse medical machine-generated IFT dataset by using GPT-4 and ChatGPT for model tuning to better align the model with various medical user intents. Our follow-up works (Xie et al., 2024; Tran et al., 2023) consistently show the benefits of tuning medical LLMs with diverse machine-generated datasets.
@@ 3 Method
3.1 Clinician-curated seed dataset
@@
A diverse and high-quality seed task is essential for prompting LLMs in automatic task generation (Wang et al., 2023b). We focus on 4 key areas, taking into account various user intents in medical applications, to improve the diversity of seed instructions: topic, view, type, and difficulty level. Specifically, the topic covers various submedical domains, such as radiology, genetics, and psychophysiology. The view is derived from diverse medical personnel, including researchers, medical students, and patients, who have different inquiries, to ensure a comprehensive range of viewpoints based on various levels of domain knowledge. For the type, we include various task formats, such as summarization, rewriting, single-hop, and multi-hop reasoning, to align with different application needs. Lastly, each task is categorized by its medical difficulty level, ranging from 1 to 5 (low to high), to ensure that the seed tasks can prompt new tasks on a wide range of expertise levels. We defer the explanation of the difficulty score to Appendix for further clarification. A clinician crafts each task considering these 4 dimensions, and each task contains instruction and may have a corresponding input, which could be a detailed medical example to further elucidate the instruction and enhance task diversity. Examples are shown in Figure 1.
3.2 Medical IFT dataset genreation and LLM tuning
@@
We utilize GPT-4 for in-context learning by randomly selecting 3 tasks from the seed set and generating 12 tasks for each run. To ensure generated task diversity, we instruct GPT-4 to consider the 4 aspects outlined in 3.1. Detailed prompt instructions are provided in Table 10 in the Appendix. To further amplify textual diversity, instructions with a Rouge-L similarity above 0.7 to any other generated task are discarded (Wang et al., 2023b). Due to the lengthy propriety of medical text, we separately generate responses for each task using ChatGPT (GPT-3.5-turbo), which has demonstrated efficacy in the medical domain (Zhang et al., 2023). Finally, we result 52k machine-generated medical instruction-response pairs, MedInstruct-52k.
@@
In IFT stage, we adopt the same training prompt and hyper-parameter setup as Taori et al. (2023) to fine-tune LLaMA models on MedInstruct-52k, Specifically, we employ instructions and inputs (when available) as inputs to tune the model to generate corresponding response outputs through a standard supervised fine-tuning with cross-entropy loss. We defer hyper-parameter setup into Appendix A.3.
4 Experimental Setup
4.1 Free-form Instruction Evaluation
Datasets. To evaluate the effectiveness of LLMs in a user-oriented manner, we conduct free-form instruction evaluations on two medical datasets. (1) iCliniq [^1] We randomly selected 1,000 instances for evaluation from the 10,000 instances proposed by Li et al. (2023b)., a dataset comprising transcripts of real patient-doctor conversations collected from an online website (Li et al., 2023b). In this task, the model processes patient inquiries as input and then simulates a doctor to provide corresponding answers. (2) MedInstruct-test, a dataset created by our clinicians, includes 216 medical instructions. These instructions mimic inquiries posed by different medical personnel, varying in difficulty on a scale from 1 to 5, with 1 being the simplest and 5 being the most challenging. We present the statistics of this dataset in Table 1, and the description of the difficulty levels is provided in the Appendix A.1.
Difficulty Level 1 2 3 4 5
Count 44 46 41 41 44
4.2 Benchmark Evaluation
4.3 Baselines
@@ 5.1 Main Results
5.2 Generalizability Evaluation
Training models with specific data may lead to catastrophic forgetting, limiting their generalizability (Kirkpatrick et al., 2017). Our approach, instruction tuning a model with a diverse, domain-specific dataset, aims to improve its generalizability simultaneously. We test this using AlpaCare in AlpaFarm (Dubois et al., 2023), MMLU (Hendrycks et al., 2021), BBH (Suzgun et al., 2022) and TruthfulQA (Lin et al., 2022). We compare AlpaCare with 4 reference LLMs in AlpaFarm and report the average score. The detailed score for each reference model is deferred to Table 13 in the Appendix. We follow Chia et al. (2023) to holistically evaluate models’ general domain knowledge on MMLU (5-shot) and BBH (3-shot), receptively; and evaluate the model truthfulness on TruthfulQA (0-shot) with Gao et al. (2021). The results are shown in Table 4.
@@ 5.2 Generalizability Evaluation
Medical LLMs typically lag behind or achieve results comparable to the general domain LLM, Alpaca, in terms of generalizability. However, AlpaCare significantly outperforms both medical and general domain baselines in multiple general tasks on average. Specifically, AlpaCare shows a significant relative improvement of 57.8% on AlpacaFarm compared to the best baseline, demonstrating strong general instruction-following ability. Moreover, AlpaCare scores higher in general knowledge tasks and maintains comparable truthfulness scores compared to other baselines, indicating strong generalization abilities due to high data diversity.
5.3 Ablation Study
To further understand the effectiveness of AlpaCare, we conduct systematic ablation studies on two medical free-form instruction evaluations and report the average score of each task across 4 reference models, receptively. The results of each reference model are detailed in the Appendix A.4.2.
AlpaCare consistently delivers superior performance in 13B model comparisons. To explore the impact of scaling up the LLM backbone, we fine-tune AlpaCare-13B on LLaMA-13B and compare its performance against other 13B LLMs.The results are shown in Table 6.
@@ 5.3 Ablation Study
Table 5: Result comparison on 13B instruction-tuned models. Comparing the average performance of AlpaCare-13B and its 13B baselines with GPT-3.5-turbo as the judge across 4 distinct reference models.
@@
Table 6: Results on different LLM backbones. Comparing the performance of Alpaca and AlpaCare using different LLM backbones.
iCliniq MedInstruct
LLaMA Alpaca 24.4 23.2
AlpaCare 53.6 53.5
LLaMA-2 Alpaca 30.3 26.8
AlpaCare 53.7 54.2
AlpaCare achieves superior performance across various backbones. To explore the effect of different LLM backbones, we tune Alpaca-LLaMA2 and AlpaCare-LLaMA2 by training LLaMA2-7B on Alpaca data and MedInstruct-52k, respectively. Table 6 compares the performance of Alpaca and AlpaCare based on different LLM backbone families.
Consistent with the results of using LLaMA-1 as the backbone, AlpaCare-LLaMA2 consistently and significantly outperforms Alpaca-LLaMA2 in both datasets. This further underscores the backbone agnostic property of our method and emphasises tuning with a diverse medical IFT dataset can bolsters models’ medical capabilities.

### v6

@@ Abstract
Instruction-finetuning (IFT) has become crucial in aligning Large Language Models (LLMs) with diverse human needs and has shown great potential in medical applications. However, previous studies mainly fine-tune LLMs on biomedical datasets with limited diversity, which often rely on benchmarks or narrow task scopes, significantly limiting the effectiveness of their medical instruction-following ability and generalizability. To bridge this gap, we propose creating a diverse, machine-generated medical IFT dataset, MedInstruct-52k, using GPT-4 and ChatGPT with a high-quality expert-curated seed set. We then fine-tune LLaMA-series models on the dataset to develop AlpaCare. Despite using a smaller domain-specific dataset than previous medical LLMs, AlpaCare not only demonstrates superior performance on medical applications, with up to 38.1% absolute gain over best baselines in medical free-form instruction evaluations, but also achieves 6.7% absolute gains averaged over multiple general domain benchmarks. Human evaluation further shows that AlpaCare consistently outperforms best baselines in terms of both correctness and helpfulness. We offer data, model, and code publicly available  [^1] https://github.com/XZhang97666/AlpaCare.
@@ 1 Introduction
Recent advancements in the training of large language models (LLMs) have placed a significant emphasis on instruction-finetuning (IFT), a critical step in enabling pre-trained LLMs to effectively follow instructions (Ouyang et al., 2022; Longpre et al., 2023; Wei et al., 2022). However, relying solely on NLP benchmarks to create instructional datasets can lead to ‘game-the-metric’ issues, often failing to meet actual user needs (Ouyang et al., 2022). To better align with human intent, Wang et al. (2023b) introduces the concept of fine-tuning LLMs using diverse machine-generated instruction-response pairs. Subsequent works further highlight the importance of diversity in IFT datasets (Taori et al., 2023; Xu et al., 2023a; Chiang et al., 2023; Chen et al., 2023). However, how to improve dataset diversity in the medical domain for aligning with various user inquiries is still underexplored.
LLMs have demonstrated significant potential in the medical domain across various applications (Nori et al., 2023a; b; Singhal et al., 2022; 2023; Liévin et al., 2023; Zhang et al., 2023). To alleviate privacy concerns and manage costs, several medical open-source LLMs (Han et al., 2023; Li et al., 2023b; Wu et al., 2023; Xu et al., 2023b) have been developed by tuning LLaMA (Touvron et al., 2023a; b) on medical datasets. Even substantial volumes, these datasets are limited in task scopes and instructions, primarily focusing on medical benchmarks or specific topics, due to the high cost of collecting real-world instruction datasets (Wang et al., 2023b), particularly when extending further into the medical domain(Jin et al., 2021; 2019). This lack of diversity can negatively impact the models’ ability to follow instructions in medical applications and their effectiveness in the general domain. Therefore, there is an urgent need for a method to generate diverse medical IFT datasets that align with various domain-specific user inquiries while balancing cost.
@@
To automatically generate a broader array of tasks for training, we prompt GPT-4 to create instructions for new medical tasks by leveraging the existing clinician-curated tasks as demonstrations. After generating tasks and conducting deduplications, we employ ChatGPT to provide responses to the valid tasks. Consequently, we compile a 52k medical self-instruct dataset, MedInstruct-52k, which supervises the tuning on the LLaMA series models (Touvron et al., 2023a; b; AI@Meta, 2024), resulting in AlpaCare. Due to the limited number of test sets available for evaluating medical LLMs in terms of instruction-following ability and medical capacity, we introduce a new clinician-crafted free-form instruction evaluation test set, MedInstruct-test, covering medical tasks across different difficulty levels.
@@
  • We propose a semi-automated, diverse medical IFT dataset generation pipeline to produce cost-effective, high-quality IFT data for LLM alignment in various medical applications.
@@
  • We release MedInstruct-52K, a diverse machine-generated medical IFT dataset with 52K instruction-response pairs, and MedInstruct-test, a test set of 216 clinician-crafted medical tasks, to help build and evaluate medical LLMs.
@@ 2 Related Works
IFT. Closed-form IFT creates IFT datasets from existing NLP benchmarks using carefully designed instructions to improve model generalization on new tasks (Wei et al., 2022; Sanh et al., 2022; Chung et al., 2022; Longpre et al., 2023). However, these instructions are often simpler than real-world scenarios, leading to models that fail to align with diverse user intentions. In contrast, Ouyang et al. (2022) collects a diverse IFT dataset with real-world instructions and responses, rich in both instruction forms and task types. They train GPT-3 Brown et al. (2020) on this dataset to obtain InstructGPT (Ouyang et al., 2022), demonstrating promising results in aligning with diverse actual user needs. Due to the closed-source propriety of strong LLMs (e.g. ChatGPT and GPT-4), various open-source instruction fine-tuned models (Taori et al., 2023; Xu et al., 2023a; Chiang et al., 2023; Peng et al., 2023) have been proposed to tune open-source LLMs using datasets obtained from these strong teacher models to enhance their instruction-following abilities. Alpaca (Taori et al., 2023) creates a 52k diverse machine-generated IFT dataset by distilling knowledge from the "teacher" Text-Davinci-003 (Hinton et al., 2015; Li et al., 2022). Peng et al. (2023) utilizes the same instructions with Alpaca but adopts GPT-4 as the "teacher" LLM to generate higher-quality and more diverse responses to improve the model’s alignment on 3H (Helpfulness, Honesty, and Harmlessness) (Askell et al., 2021). Vicuna (Chiang et al., 2023) is trained on the ShareGPT data (Sharegpt, 2023), which contains actual ChatGPT users’ diverse instructions, obtaining strong response quality and instruction-following ability. However, creating diverse IFT datasets for aligning models with various user intentions in the medical domain remains underexplored.
@@
LLMs in Biomedicine. Closed-source LLMs have demonstrated significant proficiency in the medical domain (OpenAI, 2022; 2023; Singhal et al., 2023; 2022). ChatGPT demonstrates promise in the US Medical Exam (Kung et al., 2023) and serves as a knowledge base for medical decision-making (Zhang et al., 2023). The MedPaLM (Singhal et al., 2022; 2023) have shown performance in answering medical questions on par with that of medical professionals. GPT-4 (OpenAI, 2023) obtains strong medical capacities without specialized training strategies in the medical domain or engineering for solving clinical tasks Nori et al. (2023a; b). Due to privacy concerns and high costs, several open-source medical LLMs (Xu et al., 2023b; Han et al., 2023; Li et al., 2023b; Wu et al., 2023) have been built by tuning open-source base models on medical corpus. ChatDoctor (Li et al., 2023b) is fine-tuned using 100k online doctor-patient dialogues, while Baize-Healthcare (Xu et al., 2023b) employs about 100k Quora and MedQuAD dialogues. MedAlpaca (Han et al., 2023) utilizes a 230k dataset of question-answer pairs and dialogues. PMC-LLAMA (Wu et al., 2023) continually trains LLaMA with millions of medical textbooks and papers, and then tunes it with a 202M-token dataset formed by benchmarks and dialogues during IFT. However, due to the high cost of collecting diverse real-world user instructions (Wang et al., 2023b), their datasets are limited in diversity, mainly focusing on medical benchmarks or within certain topics, such as doctor-patient conversations, hampering models’ medical instruction-following ability and generalizability. We propose creating a cost-effective, diverse medical machine-generated IFT dataset using GPT-4 and ChatGPT to better align the model with various medical user intents. Other follow-up works after ours (Xie et al., 2024; Tran et al., 2023) consistently show the benefits of tuning medical LLMs with diverse machine-generated datasets.
@@ 3 Method
Clinician-curated Seed Dataset
@@
A diverse and high-quality seed task is essential for prompting LLMs in task generation (Wang et al., 2023b). We focus on 4 key areas, taking into account various user intents in medical applications, to improve the diversity of seed instructions: topic, view, type, and difficulty level. Specifically, the topic covers various submedical domains, such as radiology, genetics, and psychophysiology. The view is derived from diverse medical personnel, including researchers, medical students, and patients, who have different inquiries, to ensure a comprehensive range of viewpoints based on various levels of domain knowledge. For the type, we include various task formats, such as summarization, rewriting, single-hop, and multi-hop reasoning, to align with different application needs. Lastly, each task is categorized by its medical difficulty level, ranging from 1 to 5 (low to high), to ensure that the seed tasks can prompt new tasks on a wide range of expertise levels. We defer the explanation of the difficulty score to Appendix B for further clarification. A clinician crafts each task considering these 4 dimensions, and each task contains instruction and may have a corresponding input, which could be a detailed medical example to further elucidate the instruction and enhance task diversity. Examples are shown in Figure 1.
Medical IFT Dataset Generation and LLM Tuning
@@
We utilize GPT-4 for in-context learning by randomly selecting 3 tasks from the seed set and generating 12 tasks for each run. To ensure generated task diversity, we instruct GPT-4 to consider the 4 aspects outlined in 3. To further amplify textual diversity, instructions with a Rouge-L similarity above 0.7 to any other generated task are discarded (Wang et al., 2023b). Due to the lengthy propriety of medical text, we separately generate responses for each task using ChatGPT (GPT-3.5-turbo), which has demonstrated efficacy in the medical domain (Zhang et al., 2023). Finally, we result in 52k machine-generated medical instruction-response pairs, MedInstruct-52k. Detailed prompts for instruction and output generation are provided in the Appendix C. To verify medInstcut-52k’s data quality, we randomly select 50 instances for a clinician to evaluate, resulting in 49 out of 50 responses being graded as correct, which demonstrates the dataset’s high quality. The LLM APIs cost analysis and chosen reason are in Appendix D.
@@
During IFT, we adopt the same training prompt and hyper-parameter setup as Taori et al. (2023) to fine-tune LLaMA models on MedInstruct-52k, Specifically, we employ instructions and inputs (when available) as inputs to tune the model to generate corresponding response outputs through a standard supervised fine-tuning with cross-entropy loss. We defer hyper-parameter setup into Appendix E.
@@
4 Experimental Setup
Free-form Instruction Evaluation
Datasets.To evaluate the effectiveness of LLMs in a user-oriented manner, we conduct free-form instruction evaluations on two medical datasets. (1) iCliniq [^2] We randomly selected 1,000 instances for evaluation from the 10,000 instances proposed by Li et al. (2023b)., a dataset comprising transcripts of real patient-doctor conversations collected from an online website (Li et al., 2023b). In this task, the model processes patient inquiries as input and then simulates a doctor to provide corresponding answers. (2) MedInstruct-test, a dataset created by our clinicians, includes 216 medical instructions. These instructions mimic inquiries posed by different medical personnel, varying in difficulty on a scale from 1 to 5, with 1 being the simplest and 5 being the most challenging. We defer the difficulty levels description and test set statistics into the Appendix B and F, respectively.
Benchmark Evaluation
Baselines
5 Experiment Results
Main Results
@@ 5 Experiment Results
Generalizability Evaluation
Training models with specific data may lead to catastrophic forgetting, limiting their generalizability (Kirkpatrick et al., 2017). Our approach, instruction tuning a model with a diverse, domain-specific dataset, aims to improve its generalizability simultaneously. We test this using AlpaCare in AlpaFarm (Dubois et al., 2023), MMLU (Hendrycks et al., 2021), BBH (Suzgun et al., 2022) and TruthfulQA (Lin et al., 2022). We compare AlpaCare with 4 reference LLMs in AlpaFarm and report the average score, and follow Chia et al. (2023) to holistically evaluate models’ general domain knowledge on MMLU (5-shot) and BBH (3-shot), receptively; and evaluate the model truthfulness on TruthfulQA (0-shot) with Gao et al. (2021). The results are shown in Table 3. The detailed score for each reference model on AlpaFarm and more general domain experiment are deferred to Table 13 and Table 14 in the Appendix G.
Medical LLMs often have worse or comparable results than the general LLM, Alpaca, in terms of generalizability. However, AlpaCare significantly outperforms both medical and general domain baselines in multiple general tasks on average. Specifically, AlpaCare shows a significant relative improvement of 57.8% on AlpacaFarm compared to the best baseline, demonstrating strong general instruction-following ability. Moreover, AlpaCare scores higher in general knowledge tasks and maintains comparable truthfulness scores compared to other baselines, indicating strong generalization abilities due to high data diversity.
Ablation Study
To further understand the effectiveness of AlpaCare, we conduct systematic ablation studies on two medical free-form instruction evaluations and report mean results of each task across 4 reference models, receptively. We defer detailed results of each reference model into Appendix G.
AlpaCare consistently delivers superior performance in 13B model comparisons. To explore the impact of scaling up LLM size, we fine-tune AlpaCare-13B on LLaMA-13B and compare its performance against other 13B LLMs. Results are shown in Table 5.
@@
Table 4: Result comparison on 13B instruction-tuned models.
@@
Table 5: Results on different LLM backbones.
LLaMA LLaMA-2 LLaMA-3
iCliniq Alpaca 24.4 30.3 26.8
AlpaCare 53.6 53.7 56.9
MedInstruct Alpaca 23.2 26.8 20.7
AlpaCare 53.5 54.2 56.6
AlpaCare achieves superior performance across various backbones. To explore the effect of different LLM backbones, we tune Alpaca-LLaMA2/3 and AlpaCare-LLaMA2/3 by training LLaMA2-7B (Touvron et al., 2023b) and LLaMA3-8B (AI@Meta, 2024) on Alpaca data and MedInstruct-52k, respectively. Table 5 compares the performance of Alpaca and AlpaCare based on different LLM backbone families.
Consistent with the results of using LLaMA-1 as the backbone, AlpaCare-LLaMA2/3 consistently and significantly outperforms Alpaca-LLaMA2/3 in both datasets. This further underscores the backbone agnostic property of our method and emphasises tuning with a diverse medical IFT dataset can bolsters models’ medical capabilities.